Astronomy Report – January 24th

Crew Astronomer: Santiago Vargas Dominguez


Missing exposure for Horsehead Nebula (Barnard 33) was completed and I processed the complete set of nearly 3 hours in LRGB+Halpha with multiple exposure times (50, 120 and 300 seconds). Please find final cropped image attached.

Objects Viewed: Barnard 33.

Problems Encountered: Dome was close due to the problems with the generator

Journalist Report – January 24th

Tittle: The Great Journey.

Author:Hermes Bolivar.

Sun number 12 of the mission, The Sun rises and each of the crew are available to their daily tasks and at the same time begins the expectation of their return to Earth, in each of us rolls the uncertainty of our return, everything that awaits in our lives, return to the routines of work and research, to cities full of people, with their daily problems.

Mars has offered us a hostile environment, but in turn unimaginable landscapes and a tranquility that has no price, that is what undoubtedly those crew members who left today to their extra vehicular activity thought. This type of environments that are increasingly less frequent on Earth stimulate our imagination and makes us reflect on how small we are in this vast and mysterious universe. It also makes us understand that this new world is just a small step through this great journey that we have started from our origin as a species.

Meanwhile, today I had to stay at the station, taking care of its maintenance and research projects. Some of them without problems, others of them that still require our attention. In the GreenHab, my job, the plants with their green color and their variety of aromas remind me of those mountainous fields of the Colombian Andes from which I come and miss.

Only a few days after returning to Earth, there are many lessons that I have from this place. I have understood that before this adventure comes to its conclusion, the experience we have had here is like the seed of one of the many plants that I have sown in the GreenHab, and that from this, the generation that will start the Great journey through the cosmos shall come.

Journalist report

Titulo: El Gran Viaje.

Autor: Hermes Bolívar

Sol numero 12 de misión, Amanece y cada uno de los tripulantes se disponen a sus quehaceres diarios y al mismo tiempo comienza la expectativa de su vuelta a la tierra, en cada uno de nosotros rueda la incertidumbre de su regreso, de todo lo que nos espera en nuestras vidas, volver a las rutinas del trabajo y de la investigación, a las ciudades llenas de personas, con sus problemas cotidianos.

Marte nos ha ofrecido un ambiente hostil, pero a su vez paisajes inimaginables y una tranquilidad que no tiene precio, eso es lo que sin duda pensaron aquellos tripulantes que salieron el día de hoy a su actividad extra vehicular. Este tipo de ambientes que cada vez son menos frecuentes en la tierra estimulan nuestra imaginación y hace que reflexionemos acerca de lo pequeños que somos en este basto y misterioso universo. También nos hace entender, que este nuevo mundo es apenas un pequeño paso a través de este gran viaje que hemos iniciado desde nuestro origen como especie.

Mientras tanto, hoy yo tuve que quedarme en la estación ocupándome de su mantenimiento y de los proyectos de investigación. Algunos de ellos sin inconvenientes, otros de ellos que aun requieren de nuestra atención. En el GreenHab, mi puesto de trabajo, las plantas con su color verde y su variedad de aromas me hacen recordar aquellos campos montañosos de los andes colombianos de los cuales vengo y extraño.

A solo pocos días de regresar a la tierra, son muchas las lecciones que me quedan de este lugar. He comprendido que antes de que esta aventura llegue a su conclusión, la experiencia que hemos tenido aquí es como la semilla de una de las muchas plantas que he sembrado en el GreenHab, y que de esta, germinara la generación que dará el inicio al gran viaje a través del cosmos.

GreenHab Report – January 24th

[category GreenHab-report]

Greenhab Officer:

Hermes Bolivar

Environmental Control:


Cooling w/ ambient air (7 hrs)

40% Shade Cloth on

80% Shade Cloth on

Average Temperatures:

Low: 15.8°C

High: 26.4°C

Hours of Supplemental Light:


Daily Water Usage of Crops:

15 gallons

Water in Blue Tank (gallons):

39.3%, 118 Gallons.

Times of Watering for Crops:






Changes to crops:

* we plant two trays of microgreeens and babygreens.


* I plant microgreens and baby greens on two trays available, with the help of two crewmembers. The greeenhab won’t have supplemental light because the energy generator doesn’t work.


* 4 g of parsley.

Support/Supplies Needed:


Operations Report – January 24th

SOL: 12

Name of person filing report: Freddy Castaneda

Non-nominal systems: Nothing new to report.

Notes on non-nominal systems: None

Generator (hours run): 7hr 0min; Turned on last night (23Jan2019) at 16:30; Stopped running during the night, about 23:30. It was turned off all night. This morning the contractor checked the generator, it will be turned off for three days.

Solar SOC – Turned on (23Jan2019) 84%; Turned off (24Jan2019) 95%; Turned on (24Jan2019) 86%

Diesel Reading – Under 1/2 tank.

Propane Reading – Still 80%

Ethanol Free Gasoline – Nothing to report.

Water (auxiliary tank) – Not in use.

Water (static tank) – Just over 20%; 115 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: Not applicable

Water in GreenHab – About 37%; 110 gallons

Water (loft) – At level marker 12

Static to Loft Pump used – Yes; At 18:05 to refill tank

Water Meter: 01404464

Toilet tank emptied: No.

Deimos rover used: No, still not functional

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Not used.

Hours: 70.3

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Opportunity rover used: Not used.

Hours: 46.2

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Not used.

Hours: 72.2

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: Brakes on Opportunity still not working properly.

ATV’s Used: None (Honda, 300, 350.1, 350.2, 350.3)

Reason for use: None

Oil Added? No

ATV Fuel Used: None

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? Not used.

CrewCar used and why, where? Off campus

General notes and comments: Nothing to report.

Summary of internet: Nothing to report.

Summary of suits and radios: Suit #3 without fuse, Suit #10 Pending to check battery functionality on field, Suit #11 Pending to check battery functionality on field.

Summary of Hab operations: Vacuum filter and tissue paper received.

Summary of GreenHab operations: Nothing to report.

Summary of ScienceDome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: Nothing to report.

EVA Report – January 24th

EVA # 8

Author: Liza Forero

Purpose of EVA:

A walk to make a regional geologic recognition of the western part of the MDRS zone.

Start time:

End time:


The purpose of this EVA was to make a regional recognition of the geologic features of the zone. We went by walking to the Hab Ridge and we decided not to go to Pooh´s corner so we didn´t make use of the rovers this time.

The weather was okay and we did not have any trouble with suits.

Destination: Hab Ridge (walking)

Coordinates: 517600 E, 425000 N

Oscar Ojeda, Freddy Castañeda, Yael Mendez, Liza Forero

Road(s) and routes per MDRS Map: We surrounded the HAB and walk to west to get to the Hab Ridge

Mode of travel: Walking

Vehicles used: None

Sol Summary – January 24th

Sol: 12

Summary Title: Europa VII

Author’s name: Oscar Ojeda

Mission Status: Nominal

Sol Activity Summary: Today was a fine day, we have already spent 12 days on Mars, and although we are exhausted, laughter and good mood is everywhere (The auto-corrector suggests me to change it to "Good Food", which is also true). As always, we started the day with a big breakfast, those powdered eggs are starting to taste good, and supplies will last perfectly until our last meal on Mars. We got ready for our last EVA of the mission, and while we traversed the hab ridge, the crewmembers on the station were working on the final touches of their research. We came back to a plentiful lunch, and the afternoon was spent in more research. We expect to watch the sky tonight, as we get ready to say goodbye to the pale red dot, our home, our present, and hopefully, our future.

Look Ahead Plan: We have two days left here on Mars. After almost two wonderful and demanding weeks, we’re almost ready to come back to Earth. There are some final results to take from the research, finalize the global reports, and bring the mission to a close.

Anomalies in work: None

Weather: Clear skies all day long, low to moderate intensity winds.

Crew Physical Status: In good health, in better shape than yesterday…!

EVA: 1, a team of 4 went to the Hab Ridge.

Reports to be filed: Sol Summary, Operations Report, Greenhab Report, Journalist Report, EVA Report, HSO Report, Astronomy Report.

Support Requested: Can you send us some more news from Colombia?

Copyright © The Mars Society. All rights reserved. | Main Site